From GlueXWiki
Jump to: navigation, search

June 12, 2014 FDC+CDC meeting


  1. Gas system
  2. DAQ (Beni)
  3. CDC status (Mike)
  4. Engineering (Bill)
    • new cathode measurements [1] (Lubomir)
  5. Electronics (Chris, Nick)
  6. Readiness review (Beni, Lubomir)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis (CMU), Bill, Beni, Dave, Eugene, Chris, Nick, Simon, Sergey Furletov, Fernando, and Lubomir (JLab).

Gas system

- The new gas control box has arrived and yesterday Scott replaced the faulty one at the FDC mixing tank. When doing this Scott was following the settings in the procedure where the channels for CO2 and Ar are swapped. However, as it turned out after the meeting, he also followed the cables on the back of the box and swapped them too, in result the gas mixture didn't change (there was some confusion during the meeting).

- Several times during the past week the HV tripped both in FDC and CDC, and we think it could be correlated with the change of the atmospheric pressure. We need to measure, record and archive the absolute pressure since the gas properties scale as E/p (filed/pressure). Dave will look for solution.

- The rates of filling the FDC and CDC mixing tanks are very similar now but they are supposed to be 0.6 vs 0.8 l/min. We suspect some leakage, maybe in the relief valves. We are going to do some tests tomorrow with Scott to figure this out.


- Beni got the DAQ working! Bryan Moffit helped him. It is working now with 4 CDC and one FDC crate. In the FDC crate in addition to the "slow" F1TDCs, Beni has also one fast F1TDC for the trigger scintillators.

- Beni's plans: taking data with the CDC and setting the readout window, then doing the same with the FDC, then increasing the number of the crates.


- Fernando wants to make sure the temperature is OK after turning ON the LV (it was OFF so far for the DAQ tests) and that the cooling pipes are directed to the CDC. Dave proposed to install a couple of thermocouples, but where? Fernando: it is not practical to have them directly on the pre-amps (there are no cooling brackets in the CDC), but next to it to make sure the cooling is directed correctly.


- Fernando on the VME readout protocol: all the modules have been tested (single crate test) with the 2eSST protocol at 40MHz clock which gives you 160 Mby/sec. If using 80MHz it will be 320Mby/sec (max for the VME) but we don't need this (TOF crates are using 50MHz/200Mby/sec). However, in the moment the DAQ can't work with the 2eSST, instead MBLT is used at 80Mby/sec. Therefore, Fernando thinks the software is to be blamed since it worked with the single crate test software; has to be discussed with Bryan.

- Three HV1550P modules have arrived back from repair, another two had to be shipped to Italy for repair. These three modules will be installed, replacing some of the faulty/temporary CDC modules. Fernando will submit a PR for two extra 1550P spares.


- Lubomir did measurements (pictures and tables attached above) on the spare cathode using a microscope with a camera. Two scales have been used to calibrate the camera scale. One is the Fernando's precise scale, which is short and was used to measure the gaps between the strips. No big gaps (more than 60 microns above nominal) were found; the CMM measurements indicated some gaps between the strips to be ~200 microns bigger than the nominal. Absolute measurements were done with a long (6ft) metal scale. The technique is to go along the scale with the microscope, and read the positions of the strips together with the scale. Measurements were done so far only at the edges of the three foils along two lines across the strips. The results show that the pitch of the strips for the three foils is almost the same (~5.011 mm) while the CMM showed different values: between 5.003 and 5.018 mm. We suspect the CMM automatic measurement had issues at the seams between the foils due to the extra Kapton foil there.

- We certainly need more measurements with the microscope. After some discussion we came to a plan: will use the granite stringing table to put the strong-back (with the cathode) on it, and to scan it will attach the microscope to the stage. Sergey proposed to investigate another solution, to make a high resolution photo at ~5m from the cathode; we will first try to see how feasible it is.

Readiness review

- Beni and Lubomir are responsible for the CDC, FDC and gas sections. The instructions for the shift workers are ready, need to finish the checklist by the beginning of the next week.